An unknown solution containing iron at a concentration between 10 and 50 mg/L is now to be analyzed. 25 mL of this unknown solution is diluted to 250 mL with the appropriate reagents to form the ferrous - phenanthroline complex. The diluted solution has an absorption of 0.795. What is the concentration in (mg/L) of ferrous ion in the diluted solution? Enter a numeric answer only, do not include units with your answer.

part 2 ) What is the concentration of iron in the unknown solution (before it was diluted), in mg L-1 ? Enter a numeric answer only, do not include units with your answer.
Answer #1

we known that Absorbance A= ebC

where A= 0.795, e= moalr extinction coefficient of [(C2H8N2)3Fe]+2 ;  11,100 L/molcm. and b= path length= 1cm (assumed)

0.795 = 11100*1C

C= 0.795/11000=7.16 × 10-5M

atomic weight of iron =56

Mass of iron in 1 L = 7.16*10-5*56 *1000mm/L=4.010mg

moles of iron in 100ml; = 7.16*10-5*100/1000= 7.16*10-6

this many moles are there in 5ml of unknown solution

Do concentration = 7.16*10-6*1000/5 moles/L= 7.16*10-6*56*200*1000 mg/L= 80.216 mg/L

so this is not the range in which concentration is measured, assume a path length of 1.75cm

and repeat the calculations

C now = 0.795/(1.75*11000)=4.13*10-5 Moles/L

Concentration expressed in mg/L=4.13*10-5*56*1000mg/L=11.56mg/L

moles of iron in 100ml =4.13*10-5*100/1000= 4.13*10-6 moles

this many moles are there in 5ml of unknown solution

hence concentration= 4.13*10-6*1000/5= 0.000826moles/L

expressed in terms of mg/L= 0.000826*56*1000=46.256mg/L.

This is falling into the range to be measured.